a.   You are provided with the following demand and cost information for a monopolist.

P=60-0.25 Q and TC=Q2+40Q+50 Where P is price, Q is quantity of the product and TC is total cost. Determine the profit maximizing level of output and prices


1
Expert's answer
2021-06-21T14:51:58-0400

"P = 60 -0.25Q"

"TR = P.Q"

"= (60-0.25Q)Q"

"TR = 60Q - 0.25Q^2"

"MR = dTR\/dQ"

"= 60(1) - 0.5Q"

"MR = 60 - 0.5Q"

"TC = Q^2 + 40Q + 50"

"MC = dTC\/dQ"

"= 2Q + 40"

Monopolistic profit-Maximizing condition

"MC = MR"

"60-0.5Q = 2Q + 40"

"60-40 = 2Q + 0.5Q"

"20 = 2.5Q"

"Q = 20\/2.5"

"Q = 8"

a) Profit Maximizing price

Substituting into P equation

"P = 60-0.25Q"

"= 60 -0.25(8)"

= "60-2"

Profit Maximizing price "= 58"

b) Profit Maximizing level of output

"P = MC"

"60-0.25(Q) = 2Q + 40"

"60-40 = 2Q +0.25Q"

"20 = 2.25Q"

"= 20\/2.25"

Profit Maximizing level of output "= 8.89"